Today I figured out how to calculate distances between tech reps on the NMDS plot to numerically validate my removal of poor-quality reps. I ended up removing a few more reps (as compared to visually inspecting reps), but as a whole not much has changed. I also generated a couple plots using Plotly, which is fantastic. Plotly creates interactive plots so you can hover over points, zoom into a plot, etc.
It’s been 1 week since I moved my Oly seed to the dock; today I checked on them to ensure the screen envelopes are still secured and to rinse them with fresh water. Everything was still in place and the screen wasn’t too dirty, so I’ll wait ~10 days to 2 weeks to return. I also tagged the Oly cages 92. NOTE: there are three cages hanging together on 92; my Oly broodstock and some of Yaamini’s gigas are in 2 cages, and my Oly seed are in the other.
My post-set Oly’s have been housed in upwelling silos in a tank at Manchester, fed algae produced by PSRF, since July. It’s time to get them out of there, since PSRF is really only producing algae for me, and they have potential plans to turn the water off for re-plumbing projects, etc. My task for today is to move the oysters to cages hanging off the dock.
Check out my recent Jupyter Notebook entry where I perform a full SRM Analysis.
This is vast. Lots to do.
The following is data on Oly survival from larvae to juvenile. Data is adjusted for a few events where larvae were lost due to overflowed bucket, etc. Data was adjusted by calculating the mean % larvae accounted for from one bucket count to the next (pretty good, 94%), then taking the difference between the actual # counted during screening after loss event and what would be predicted by the 94% accountability rate.
I’ve been wrastling with with R to analyze my SRM data. Here’s a Jupyter notebook with initial analysis: SRM Analysis Jupyter Notebook
One of the first steps in processing SRM data is to confirm that the selected peaks actually represent the peptides, aka that our assay works. To do this, we use linear regression between PRTC retention times in DIA and SRM to calculate predicted transition RTs in collected SRM data. Then, we calculate the R^2 for PRTC and experimental peptides compared to predicted.
Before I export my data from Skyline for data analysis, I have the following final things to do:
- Adjust peak boundaries as necessary
- Remove transitions that do not align with the predicted RT. Remove any peaks that do not have at least 2 transitions.
Screened all oysters up to 450um screen, counted set on 180um silos
First steps, done on 8/8
Here’s a quick visualization of the Olympia oyster spawning data from this spring. Each chart represents a population, and within the charts data is color coded by treatment.
Prepping dilution curve samples
How many samples can I run?
@ 8:30am on Monday 7/24 I got my samples started again. I need to figure out how many more samples I can run within the given timeline. Here are the considerations:
- Running samples in batches of (5 samples) + (1 QC) + (1 blank) = 500 minutes total, so that’s 100 minutes / sample.
- We have until 10am on Friday 7/28
- Need 40 hrs for the dilution curve run
- So, need to be done with my samples by Wednesday 7/26 @ 6:00pm
- Time between this morning @ 8:30am when I re-started my samples & 7/26 @ 6pm = 57.5 hrs = 3450 minutes / 100 minutes/sample = 34.5 samples.
- I have 25 samples left to run, plus 2 blanks (Gblank & OBlank) which I could run twice.
- Should be done with my samples on Wednesday @ 5:30am, this includes 1 run of each blank.
Thoughts on whether or not to re-make samples
Up to this point I have 20 samples where a handful of peptides don’t show up from both PRTC and my samples. In PRTC there are 4/9 poor quality peptides, and in my samples 3/39 poor quality transitions. In an ideal world I would remake these ~20 samples, run the new batch twice while being careful with freeze/thaw and time out of the freezer. However, I don’t have time for 2 runs of remade samples. I could remake, then do 1 run of each in the hopes of capturing data on those 3 transitions. However, I think it’s best to have replicates of the 36 good quality transitions in my samples. Also, the 3 poor quality transitions were not in the sample proteins, so I can likely draw conclusions about protein quantification from the other 2 transitions. The differing rates of peptide degradation between samples does make me a little concerned; I’m wondering how folks take this into consideration.
Pro Tips for using UWPR mass spec
SPOILER ALERT The thought is that some of the peptides (in both PRTC and my samples) degrade quickly, and this is what’s causing the loss of signal. I loaded my samples in batches of 25, so samples sat out up to ~2.5 days prior to being injected. I’ll need to consider “time out of freezer” as a factor when analyzing my results. This is obvious in the dilution curve data, where PRTC total area shows an obvious decrease over time.
Readied last 25 peptide samples for mass spec
Stopped by Manchester for the AM to check on things.
Using DIA PRTC retention time data to predict SRM RT, then confirming that assays are present in my samples
Mass Spec refresher, final sample prep (+PRTC, transferring to autosampler vials)
DNR Geoduck Ctenidia Sample prep for SRM
Moved >450um seed to upwellers; combined 180um setters; standard maintenance water change, etc.
We are now running like a well oiled machine, employing progessive assembly practices.
Please see my Geoduck-DNR Repo, June Analyses folder for most up-to-date version of this process.
Screening day, preserved growth exp. larvae, changed growth exp water, collected new, cleaned downwellers.
Just me and Grace today, but all was very manageable.
Counted algae on broodstock manifolds, cleaned setters, SN exp. water change, new larvae, imaged growth exp.
6/28 Low impact day
- No new larvae, but NF6 Low A was “spawny”
- Water change on growth exp
- Swapped in clean banjos
- Imaged plates #1-4 from 6/26
6/29 - high impact day- screened larvae & collected new; sampled growth exp; water change, rinsed downwellers
Screened today, imaged 6/22 growth exp. & screened larvae; cleaned broodstock, yada yada yada…
Began new larval culling protcol; prepped growth exp larvae for imaging; collected new larvae; water change on growth exp.
Looked at downwellers for set, fixed growth exp. larvae from yesterday
Setters get bigger home, sampled SN growth trial, screened larvae
Trust but verify, be explicit
Screening day and more! Things went very smoothly today. Grace, Olivia and I make a good team.
Screening day- Arrived at 8:30am, geared up to screen through the larvae. Grace worked the screening table, Olivia managed the buckets and whatnot, I counted and re-stocked larvae, and Katie took video and images. Finished at ~2:00pm. Things went very smoothly and we got a good routine down. Here are snap-shots of the data:
Maintenance day. Steven and Beyer helped me, and it was a productive morning.
Screening day, working solo.
Tasks for the day:
- Upwelling tank:
- Got water flowing in the upwelling tank outside, cleaned and installed connections, collected and tested hanging silos.
- Installed immersion heater and monitored over the day
- Cleaned setting tanks
- Inspected larvae/culch in a silo by screening through 450 onto 100; viewed under microscope for setters. Did not see any, but larvae were still present.
- Did not screen/inspect all silos; instead, cleaned all silos, drained and cleaned setting tanks, replaced drippers & air stones.
- Installed outflow tubing for clean drainage.
- Build freshwater adapter for manifolds. Need to run freshwater through system ASAP.
- Collected new larvae:
- HL-6 Low
- K-6 Low
- K-10 Ambient
- Cleaned all K & HL broodstock fully, here’s the configuration before cleaning:
The afternoon view:
Big Screening Day. Had help from Grace, Steven, and Yaamini. Screened all larval buckets through 224 & 180um, and caught the rest on 100um. Larvae that held on 224um screen were moved to the downwelling setting tanks with microcultch at 224um. The SN & NF groups were split into two buckets, 180um and 100um (aka 180um < x <224um, and 100um < X < 180um). I did not split the HL & K groups (limited on materials and space).
From where I stand, these days:
May 29th, 2017
Here are some charts of the Oly larvae that I’ve collected & counted to date
Geoduck sample selection for next round of MS/MS
Arrived @ 9:30am
- lab meeting
- can steal 2 HOBOs from gigas, leave 1
- Brent suggested that it will take ~24 hrs of dual-bucket system to successfully collect only live larvae. This is a good idea, however I don’t expect there to be sufficient room - but maybe, if larval collections are staggered… ?
- NEED TO: purchase more 5-gal buckets!
- Imaged larvae collected on 5/17, and also imaged larvae collected and saved over past week.
- Images include triplicate samples from each collection/group (see larval collection data sheet): 1x, 2x, 3x of each on Nikon SMZ645
- Will save to GoogleDrive, then to GitHub
- Checked catchment buckets
- SN-10 amb B: little bit
- NF-10 amb B: teensy amount
- SN-10 low B: a bit
- SN-6 amb B: little bit
- K-10 low: some!
- HL-10 low: maybe, probably just spermy though
- All except HL-10 have spawned within the past 4 days, and since there is no significant amount of larvae here I will not keep them- I have a suspicion that these are from the same fertilization event, and they are trickling out of the female. Will collect for real tomorrow…
- Stole 2 HOBOs from gigas (left 1 there); downloaded the data from the HOBOs and emailed to Yaamini, then reprogrammed and launched to record T every 15 minutes, installed in broodstock buckets.
- Counted larvae collected on 5/17 & imaged larvae for size analysis later
- PSRF has agreed to collect/sample my larvae tomorrow (Friday, 5/19) while I’m on campus. I pre-labeled larval rearing buckets, got water flowing, and left instructions with Jade & Alice.
- Met with Alice to get a summary of her larval husbandry practices. PSRF hasn’t updated theirs in a couple years, and there are many lessons learned. See write-up in my GitHub (I will update as needed).
- Screened and counted larvae, imaged. Larvae went down drain today.
- Temperature is holding around 17.5
It’s been 14 full days since we moved broodstock to their separate “chambers.” From the literature, oysters release larvae on average 10-12 days after fertilization. Today I will clean all broodstock, larval catchment buckets, after which I will plan to collect larvae to save/rear.
Following up on my Proteins of Interest, Part II analysis, I am taking a second, simpler look at the peak area data. This time, I am simply taking the total peak area for each protein and averaging across each treatment:
To determine over/under-expressed proteins eelgrass vs. bare treatments I did the following:
Ups and downs this weekend!
With demultiplexed files in Skyline I can export my results to .csv file for analysis. While I do still need to create a Retention Time Calculator and apply to data in Skyline, I’m taking an initial stab at finding differentially expressed proteins.
Olys needed a good cleaning today.
Which samples to extract next?
Checked on the spawning setup today and made some minor modifications.
Thanks to Steven & Doug for getting the Avtech probes online so we can 1) keep an eye on temp and pH remotely, and 2) download data!
For the past few weeks Olys were all housed in a 100L tank connected to a heater/chiller, which we used to increase the temp to promote gametogenesis. The goal was 1 degC per day up to 18degC, but we weren’t able to maintain that high temp due to a weak heater, and trying to keep temp high by reducing the flow rate caused the pH to drop, so that didn’t work. I’ll download HOBO temp data, but the Oly’s were kept around 14degC from 4/11 -> 5/2. This is what the setup looked like:
I’m currently importing my demultiplexed Lumos files into Skyline for the Geoduck DNR outplant study, using a .blib file that Emma generated via Pecan with all files, and the Geoduck gonad transcriptome as the database (database also has PRTC protein). Here are some screen shots of the peaks during import!
Emma ran Pecan with my geoduck samples and produced a .blib file; the next step is to run the .blib and my sample files through Skyline.
Organized the HOBO temperature data, and pulled some plots to get an initial glimpse
Just a quick update on some plumbing that needs to be done prior to getting my animals in buckets to spawn:
I prepped the histology samples taken on 4/8 and 4/13 to be sent off for slide preparation. My samples are Ostrea lurida whole visceral mass, with the goal of analyzing gonad maturation, and Yaamini’s samples are exclusively Crassostrea gigas gonad. I sampled NF, SN & HL populations on 4/8 and K populations on 4/13. Megan, Grace, Kaitlin, and Rhonda helped me by sampling the NF, SN & HL groups, and I did the K groups myself.
I’m in the midst of my first Pub-a-thon, a Roberts Lab competition. Thus far I’ve drafted methods & some background/intro language that I’ve pulled from my Oly proposal. Things have stagnated recently; it’s time to get serious. Here are the 14-day goals:
Let’s revisit the DNR geoduck mass spec data!
April 13th, 2017: Returned to Manchester solo to sample the K population (Katherine’s oysters spawned in 2015)
The next stage of the experiment is to condition my Olys in preparation for spawning. To do so, I need to get them on a system where I could gradually (1degC/day) raise the temperature of their tank from ambient (~10degC) to 18degC. Worked with Ryan Crim to figure out how best to do this:
- Condensed all Olys into one 100L tank.
- Sourced water from the ambient line that is not connected to the OA system; this line is filtered down to 5um (upon entering the hatchery). Set flow rate to 1.5 L/min.
- Set the algae dosing pump to 45 (this is the large pump that is installed on the wall). The water line is shared by Yaamini’s oysters and some spare oysters. All told, the flow rate on the line is ~6 L/min:
- Connected tank to a pump that recirculated water through the Teco TANK TK-500 1/6 HP Aquarium Chiller (also heats!).
- Set heater to 11degC (51.8degF). Instructed PSRF to increase temp to 12degC tomorrow.
April 8th, 2017: Lots of sampling today with help from the Ladies of the Roberts Lab
Megan, Rhonda, Grace, Kaitlyn, Yaamini, and myself (shoot, we forgot to take a photo, but here’s us feeding my chickens oysters. We also terminated the OA stage of this project.
Good news. We’ve actively used the OA system for 50 days with no major catastrophes. Celebrated by spending the day water sampling and cleaning oyster poop; here’s the summary:
Let’s review March goals to see how I did:
Upon arrival, pH & temp looked good, all tanks were flowing.
Yaamini & I spent the day at the hatchery cleaning and taking water chemistry. Here’s the report.
Quick morning stop @ Manchester to tend the oysters. Here’s the report.
I restarted Pecan using 1 of my geoduck .mzML files, and the full digested c. gigas proteome.
It was simple to start, and notice that I was successfully able to use the
--backgroundProteome input! The concern is, as always, whether Emu has enough memory to complete all 80 isolation scheme windows for this one file. Here’s a summary:
Full day at the hatchery working solo; lots accomplished! No major surprises, except for one Pacific oyster mortality in the same tank as last week; it looked like the same oyster that appeared sickly, so while I’m not surprised, 2 morts in one tank is concerning. We will start draining/vortexing weekly, rather than bi-weekly.
Final exam today, so Yaamini, Grace & Olivia did the daily maintenance at Manchester. Here’s what went down.
2/22-3/10: Pecan ran for nearly 3 weeks, and although it appeared to have been functioning correctly Sean discovered that there was a problem: not enough memory to save all the feature files (there should be 80 per sample; 1 per isolation window). It would simply move on to the next sample, and thus I wasn’t getting all the peptides analyzed. Check out Sean’s notebook entry for more details.
Pecan has been running since February 22nd. That’s almost 2 weeks! After several days of trying to log on to Emu to check in on the progress, I finally was able to get some information and view logs:
It’s been 21 days since the oysters started their OA treatments, and we spent the afternoon doing the standard weekly maintenance. Things are going smoothly! So far no major catastrophes, and water conditions are holding relatively steady.
Check out Yaamini’s post for details, since I wasn’t there!
Pecan! Lesson of the month: Proteomics bioinformatics is very intense.
It’s been 2 week’s since the OA experiment started, so today was an intensive cleaning and re-organizing day. But first, Yaamini and I spent the morning at the monthly hatchery meeting, and learning about how Olympia oysters swim around in OA from Western Washigton crew (Shawn Arellano and her students). Then…
Here’s my laundry list. Obviously, I like bullets:
Here’s a quick look at the continuous temperature data. Notice the tight cycling in the “cold” treatment due to the chiller, and the slower, diurnal cycling in the “ambient” (aka warm!) group:
We check on our oysters every Monday and Wednesday. Mondays are minimum effort days, where we clean filters, collect water chemistry samples and measurements, clean algae lines with bleach, and make sure everything is flowing and pH/T are stable. Yaamini and I are dividing duties, so I’ll post links to her notebook when she goes out and I don’t.
Yaamini and I spent the afternoon with our oysters.
Popped into Manchester a couple mornings to take water chemistry.
Got the oysters into pH treatments today! This, after a couple surprises…
Spent the day @ Manchester making the final (hopefully) touches to OA system. Also sampled the F2 population (Katherine’s). Here’s a breakdown:
- Header 2, which was supposed to be ambient, was consistently reading ~7.5. This is too low, since ambient in other tanks have been reading 7.78. I disconnected the CO2 tube from the manchurian injector, and capped it. I drained the tank 1/2 way, and refilled. pH levelled off to 7.78 by the end of the day in the header (H2).
- After watching the pH in Header 1 since Wednesday, and seeing pH drop to low 6, it was obvious that adjustments were needed to the CO2 injections system. Here’s what I did:
- Drained Header 1 to 1/3 the volume, and refilled
- Lowered pressure in main CO2 line to ~15psi
- Increased the injection frequency from 120 to 180 seconds
- Decreased the injection duration from 0.8 to 0.4 seconds
- Changed set point to 7.38 <-> 7.42; since Ambient is ~7.8, I want a larger difference between treatments
- Note: I noticed water had collected again in H1’s CO2 line. I drained it again; perhaps we need a better sealant on the CO2 connection.
- It’s been ~18hrs since making htese adjustments and pH seems to be hovering around 7.4! Will continue to monitor.
- Header 1 is connected to Relay 2 for CO2 injection
- Yellow Durafet outputs to the left Honeywell screen, top, as “input pv 1”
- PInk Durafet outputs to the left Honeywell screen, bottom, as “input pv 2”
- Green Durafet ouputs to the right Honeywell screen, bottom, as “input pv 2”
- Drained all tanks, cleaned with Vortex.
- One tank has outflow drain @ bottom; capped that.
- Note: late in the day I saw 2 nicely insulated tanks from last year’s experiment stored in the Warehouse. Let’s use these next time we clean tanks * Labeled tanks: #1-3 = low pH #4-6 = ambient
- Labeled and installed HOBO temperature loggers - one logger per tank.
- Set flow rate on all tanks to 1 Liter/minute. Did not mark valves to replicate this at each cleaning - will do so next time.
Oysters are set to go into pH treatments (ambient, 7.5) on Wedneday 2/8, so on Sunday I went to Manchester to get the OA system running, so we could test the CO2 injectors and Durafets.
Full day of sampling! THANK YOU to Grace for materials prep, Megan Hintz, Olivia & Lindsay Alma for helping Yaamini and me with the sampling, and to Steven for the oversight and OA system prep.
We went out to Manchester on 1/25/2017 to check on the Olympia oysters that are already (!) undergoing temperature treatment, to assess the OA system.
Check it out! This is a screen shot of the Lumos mass spec collecting data on my geoduck samples’ peptide contents. The most important track to monitor is the top track, which is the System Pressure (psi). The long plateus correspond to the samples’ peptides being run, short plateus (I believe) correspond to cleaning between samples.
Met Emma at the UWPR in South Lake Union; she brought our samples & the acetonitrile that we used for our samples.
We transferred our samples into labeled autosampler vials. NOTE: next time we need to bring empty autosampler vials for our blanks!!!
There were a few last items that needed to happen prior to injecting samples into Mass Spec. Here they are!
January 2017 Goals
The finale: a day of desalting
Evaporating - that’s it!
Making peptides via Mini-Trypsin Digestion
Today I determined the concentration of proteins in my samples. This was done via the Thermo Scientific - 23225 - BCA PROTEIN ASSAY KIT, following the protein prep protocol and Rhonda Elliott’s recent lab notebook entry.
Today’s lab work involved homogenizing the tissue in solution, then sonicating to lyse the cells.
DNR Geoduck OA overview
After a couple months of school under my belt, here is a list of non-class related items I’d like to accomplish by December:
This is a draft research plan for a project I could execute starting winter 2016/2017.
This is a draft research plan for a project I could possibly execute winter 2016/2017. I invite any and all feedback. Thanks!
Last week I shadowed Hollie while she performed the first few steps of DNA extraction. Here’s what I learned about the process and general tips for working in that lab:
Ok, now let’s play the shell game!